OpenStudy (anonymous):

Medal :) f(x) = x + 2, find f(3). 1 6 5 −1 2. If f(x) = −4x − 2, find f(−3) 10 −14 12 −10

OpenStudy (anonymous):

You are seeing \(x\) everywhere.. Just replace that \(x\) by 3 in first case and \(-3\) in second case.. :)

OpenStudy (anonymous):

Is the first one 5

OpenStudy (anonymous):

\[f(x) = x+2 \implies f(3) = (3) + 2 = ??\]

OpenStudy (anonymous):

Yep..

OpenStudy (anonymous):

Try second on your own..

OpenStudy (anonymous):

Give a try, put \(x=-3\) there..

OpenStudy (anonymous):

12?

OpenStudy (anonymous):

Do the substitution and tell me.

OpenStudy (anonymous):

Ac ually -10?

OpenStudy (anonymous):

I mean show me.

OpenStudy (anonymous):

Sure??

OpenStudy (anonymous):

\[f(-3) = -4(-3) -2 = ??\]

OpenStudy (anonymous):

F(-3)=-4(-3)-2= 12-2=10

OpenStudy (anonymous):

Good enough to get a medal from me.. :) Good..